OpenStudy (anonymous):

Natasha rides her bike (at a constant speed) for 3 hours, helped by a wind of 3 miles per hour. Pedaling at the same rate, the trip back against the wind takes 9 hours. Find find the total round trip distance she traveled.

OpenStudy (anonymous):

Natasha rides her bike (at a constant speed) for 3 hours, helped by a wind of 2 miles per hour. Pedaling at the same rate, the trip back against the wind takes 7 hours. Find find the total round trip distance she traveled

OpenStudy (anonymous):

Let r be the bike speed in no wind. r t = d Solve the following for r: (r + 3) 3 = (r - 3) 9 r = 6 roundTripDistance = (r + 3) 3 + (r - 3) 9 Replace r with 6 in the above expression and evaluate the result. roundTripDistance = 54 miles